with what velocity should a ball be thrown vertically upwards with so that it reaches a height of 40m?

Grade:11

1 Answers

Vikas TU
6 years ago
Taking upwards positive,
From the 2nd law of eqn. we get,
0^2 = u^2 – 2*g*40
u = root(800) => 20root(2) m/s
This is the velocity to attain 40 meter distance upwards.